# A Unified Framework for N-Dimensional Visualization and Simulation: Implementation and Evaluation including 4D Boolean Operations

### v2

@@ 4 Plex: An N-Dimensional Mesh Data Exchange Ecosystem
’NORM’/’NORMD’ Required Stores the normal vectors that define the orientation of the facets in single or double precision.
’CENT’/’CENTD’ Optional Stores the cached centroid coordinates of facets to reduce computational load.
@@
The $N$-dimensional simulation environments targeted by this research require, as their foundation, an extensible data exchange ecosystem for uniformly handling high-dimensional mesh data. However, to date, no widely accepted standard format for exchanging $N$-dimensional mesh data has existed, leading to a fragmented situation where formats vary among researchers and projects. This state of affairs has been a barrier to the smooth sharing of data between different tools and research groups. To address this challenge, this study defines a systematic data exchange ecosystem for $N$-dimensional meshes, named ”Plex” (with the .plex extension), which is a file format featuring a chunk-based architecture.
@@ 4.1 Dual-Format Strategy
Research workflows that handle $N$-dimensional data present two distinct requirements: the need to inspect and verify data content in the early stages, and the need for high processing performance in large-scale computations. To meet these requirements, Plex adopts a dual-format strategy, following the design philosophy proven by the glTF format in the 3D domain.
The first format is a schema based on the standard JSON technology, which prioritizes data readability. This allows for the easy inspection and debugging of data content using a rich ecosystem of existing tools. The second is a chunk-based binary format, .plex, which prioritizes execution performance. This format is intended for fast loading by applications and efficient data storage.
The reference implementation provided in this study includes a one-way converter from the .plex format to the JSON format. This enables researchers to typically work with the high-performance binary format while exporting to a human-readable format as needed, thereby achieving a balance between execution efficiency and data transparency.
@@
The .plex format is designed with a chunk-based architecture, following the proven extensibility model of standard formats such as PNG. A file consists of a global header followed by a sequence of at least four self-descriptive chunks as shown in Table 1; the internal structure of the META chunk is detailed in Table 2. Each chunk contains a 4-character chunk type, the data length, the data payload, and a CRC32 checksum to verify data integrity. The data payload structure for the primary chunks, excluding the META chunk, is defined as follows. The payload of the FACE chunk, which defines the topology, first specifies the total number of facets as a 64-bit unsigned integer, followed by a list of vertex indices (an array of 32-bit unsigned integers) that constitute each facet. In contrast, the payload for the VERT or VERD chunks, which store vertex coordinates, contains no metadata such as counts; for a mesh with $V$ vertices in $N$ dimensions, it consists solely of a contiguous array of $V\times N$ numerical values. The respective data types are 32-bit single-precision and 64-bit double-precision floating-point numbers. The payload layout for the NORM/NRMD chunks (storing facet normals) and the optional CENT/CNTD chunks (storing centroids) also conforms to this VERT/VERD chunk format.
A parser compliant with the specification is guaranteed to safely skip any unknown chunk types it cannot interpret. This ensures that backward compatibility with existing software is maintained even if the community adds new data in the future, such as textures or animations, as custom chunks.
Furthermore, to encourage community-driven extensions, a naming convention is established that reserves identifiers containing one or more lowercase letters for use as custom chunks. This standardization will facilitate data sharing among researchers and serve as a technical foundation for improving the reproducibility of published results. Moreover, the forward compatibility guaranteed by this specification allows for future extensions, providing a sustainable infrastructure that transforms $N$-dimensional mesh data from something previously confined to individual environments into a resource that can be shared and reused by the entire community.
Figure 9: The set of base objects used in the experiments. The top row shows the 3D objects, and the bottom row shows the 4D objects. The number to the right of each object’s name is the number of facets that constitute its mesh. The 4D objects are visualized using the hyperplane slicing feature of my framework, with their 3D cross-sections at three representative $w$-coordinates (w=0.8,0,0.8)(w=-0.8,0,0.8) displayed side-by-side.
@@ 5.1 Experimental Setup
The performance evaluation in this section is not intended to be a speed comparison with existing offline methods. Instead, its purpose is to evaluate the performance scalability of the proposed CPU-based framework with respect to the scale of the processed mesh (e.g., the number of facets). All implementation and performance evaluations in this study were conducted in the following environment.
@@ 5.2 Algorithmic Performance Evaluation
Table 4: Performance evaluation results for Boolean operations. I performed union, intersection, and difference operations in 3D and 4D spaces using combinations of the base objects shown in Figure 9. For each combination, the table shows the processing time and the size (number of facets) of the resulting mesh. The processing time is the average of three trials.
@@
Table 5: Interactive performance evaluation. This table shows the rendering performance in scenes containing 1 to 20 4D objects (hypercubes with 47 facets each), while the user continuously performed movement and view rotation operations in the xzxz and xwxw planes for 30 seconds. Each value is the average of three 30 second measurement trials.
@@
5.2.1 Performance of the Direct Quickhull Implementation
As shown in Table 3, for small-scale point clouds with 50 or fewer vertices which are the intended primary use case in my framework the processing time is under 1.7 $\mathrm{ms}$ in either dimension, confirming that its performance is sufficient for this purpose.
5.2.2 Performance of N-Dimensional Boolean Operations
Next, I evaluated the performance of the $N$-dimensional Boolean operations, which are central to interactive mesh editing. I executed union, intersection, and difference operations between various objects in 3D and 4D spaces and measured their computation times. The results are presented in Table 4. In 3D space, the operations were generally completed in several tens of milliseconds, depending on the combination of objects and the type of operation. In 4D space, however, the computational load increased significantly. Even a simple combination of two Cubes required approximately 300 $\mathrm{ms}$, and in some cases, operations between two complex Random objects took over two seconds to complete. These results indicate that high-dimensional Boolean operations can be executed on a time scale of a few seconds. This processing time is considered to be within an acceptable range for an interactive workflow, where the user waits for the result of an operation before considering the next action. This demonstrates the feasibility of incorporating these operations, which were previously dominated by offline batch processing, into interactive applications.
5.3 Interactive Performance Evaluation
This section quantitatively evaluates the performance characteristics of the proposed framework. I placed multiple simple 4D objects, each with 47 facets, and measured various performance metrics as the total number of facets in the scene was increased. The experimental results are shown in Table 5. These results indicate that the performance of my framework scales with the total number of facets in the scene. Furthermore, a specific performance limit was identified: the ability to maintain 30 fps, a common benchmark for interactive operation, is limited to a total facet count of approximately 500 (equivalent to about 10 objects). This measurement provides a performance baseline for evaluating the effects of future improvements, such as a GPU implementation.
@@ 6 Conclusion and Future Work
In this paper, I proposed an interactive framework for the unified simulation and visualization of $N$-dimensional space. The core of this framework lies in the design principle of separating topology and geometry. This allowed for the integration of functionalities such as Quickhull-based mesh generation, hyperplane slicing, and $N$-dimensional Boolean operations onto a single platform. The effectiveness of this design was demonstrated through the integration of an XPBD-based physics simulation. I also proposed a control scheme for high-dimensional exploration and the Plex data exchange format.
The proposed algorithm can be extended to arbitrary dimensions; however, in this study, the evaluation was performed on the 4-dimensional implementation. Extension to five or more dimensions may introduce new performance bottlenecks due to the increased combinatorial complexity in intersection testing and tessellation. Furthermore, major future challenges include the development of a more robust $N$-dimensional non-rigid body model, application to new phenomena such as fluid simulations, and a GPU implementation of the parallelizable components currently handled by the CPU.
@@ Acknowledgments
I would like to express my sincere gratitude to Mr. Hiroaki Naito of Waseda Osaka High School for his constant and enthusiastic guidance throughout this research. I also wish to express my deepest appreciation to Professor Yasushi Homma of Waseda University for his valuable professional advice on the mathematical expressions related to Geometric Algebra used in this paper. Finally, I would like to thank the anonymous reviewers for their constructive comments, which helped to improve this manuscript.

### v3

@@ 4 Plex: An N-Dimensional Mesh Data Exchange Format
’NORM’/’NRMD’ Required Stores the normal vectors that define the orientation of the facets in single or double precision.
’CENT’/’CNTD’ Optional Stores the cached centroid coordinates of facets to reduce computational load.
@@
4.1 Background and Challenges
In the $N$-dimensional simulation environments targeted by this study, there are cases where a means for exchanging high-dimensional mesh data between projects is required. However, for $N$-dimensional meshes, there are only a limited number of common exchange formats that are widely used across the research community. In practice, project-specific formats or extensions of existing formats are often employed. This situation can lead to additional effort when sharing data across different implementations or research groups, including data conversion, validation, and interpretation of specifications.
To address this issue, this study defines a chunk-based file format, “Plex” (.plex), for the purpose of exchanging $N$-dimensional mesh data.
4.2 Relation to Existing Formats and Positioning of Plex
HDF5 and netCDF are formats designed to store multidimensional array data in a self-describing manner and can handle $N$-dimensional datasets [9, 19]. However, because their primary purpose is array storage, they do not directly specify the topology and geometry of $N$-dimensional meshes as part of their specifications.
The VTK file formats are organized as a family of formats for describing datasets that include points, cell elements, and their associated attributes [13]. XDMF adopts a model for representing data such as meshes based on XML-based metadata descriptions and references to external data [25]. In contrast, this study focuses on defining, as a single specification, the minimal elements commonly handled by mesh data parameterized by the dimension $N$. These elements are topology and geometry. This focus does not align with the design objectives of these formats.
Based on the above considerations, this study does not aim to replace existing formats; instead, Plex is positioned as a format that explicitly describes the minimal elements (topology and geometry) commonly represented in $N$-dimensional mesh data.
@@ 4.3 Dual-Format Strategy
In research workflows that handle $N$-dimensional data, two different requirements may coexist: inspection and validation of data contents at early stages, and processing performance for large-scale computations. To address these requirements, Plex adopts a design that combines a human-readable JSON format and a chunk-based binary format .plex. The human-readable format is used for content inspection and debugging, while the binary format is used to improve efficiency in loading and saving.
The present implementation includes a one-way converter from .plex to the JSON format. This enables a workflow in which data are normally stored in the binary format and exported to the human-readable format only when necessary. In this paper, this conversion is presented as a minimal means for content inspection and third-party verification of Plex data, while providing converters to VTK-family formats or HDF5 is considered out of scope and positioned as future work.
@@
The .plex format is defined as a chunk-based format consisting of a sequence of self-describing chunks. A file comprises a global header and at least four self-describing chunk sequences shown in Table 1, and the internal structure of the META chunk is shown in Table 2. Each chunk contains a four-character chunk type, a data length, a data payload, and a CRC32 checksum for verifying data integrity. The data payloads of the primary chunks, except for the META chunk, are defined as follows.
In the FACE chunk, which defines topology, the total number of facets is recorded as a 64-bit unsigned integer, followed by the vertex index lists (arrays of 32-bit unsigned integers) that constitute each facet. The VERT or VERD chunks that store vertex coordinates contain no metadata such as counts; for an $N$-dimensional mesh with $V$ vertices, they consist solely of a contiguous array of $V\times N$ numerical values. The data types are 32-bit single precision and 64-bit double precision, respectively. The data layouts of the NORM/NRMD chunks and optional CENT/CNTD chunks likewise conform to the formats of the VERT/VERD chunks.
From a compatibility perspective, parsers compliant with the specification are required to be able to skip unknown chunk types, allowing loading to continue as long as the existing required chunks can be interpreted. In addition, assuming future extensions such as additional attributes or auxiliary information to be introduced as chunks, a naming convention is defined in which identifiers containing at least one lowercase letter are reserved for custom chunks. Here, “compatibility” refers to forward compatibility within the Plex specification, whereby loading can proceed using only known chunks while ignoring unknown ones, and does not imply binary compatibility with VTK-family formats or HDF5.
Figure 9: The set of base objects used in the experiments. The top row shows the 3D objects, and the bottom row shows the 4D objects. The numerical value shown to the right of each name indicates the mesh size (number of facets). 4D objects are displayed as 3D cross-sections obtained using the hyperplane slicing function described in this paper. As an example, cross-sections at three $w$ coordinates (w=0.8,0,0.8)(w=-0.8,0,0.8) are arranged from left to right.
@@ 5.1 Experimental Setup
The performance evaluation in this chapter, for the 4D implementation, does not aim to compare execution speed with existing offline methods. Instead, it aims to evaluate the scalability of performance with respect to the size of the target meshes (e.g., the number of facets) within the proposed CPU-based architecture. All implementation and performance evaluations in this study were conducted in the following environment.
@@ 5.2 Algorithmic Performance Evaluation
Table 4: Performance evaluation results for Boolean operations. For combinations of the basic objects shown in Fig. 9, union, intersection, and difference operations were performed in three- and 4D spaces. Each entry reports the processing time averaged over three trials ($\mathrm{ms}$) and the size of the generated mesh (number of facets).
@@
Table 5: Interactive performance evaluation. Rendering performance is shown for a scene in which 1–20 4D hypercubes, each with 47 facets, are placed, while continuous interaction (translation and viewpoint rotation in the xzxz and xwxw planes) is performed for 30 seconds. As performance metrics, the average FPS, minimum FPS, and the 99th percentile frametime ($\mathrm{ms}$) are reported. All values are averages over three 30-second measurement runs.
@@
5.2.1 Convex Hull Mesh Generation Performance
The measurement results are shown in Table 3. For point sets with up to 50 vertices, which correspond to the intended input of this implementation, the processing time was less than $1.7\text{\,}\mathrm{ms}$ in all tested dimensions within the measured range.
In this implementation, convex hull generation is a processing stage whose computational cost varies depending on the input size within the architecture. To present baseline data for the architectural design regarding how convex hull generation time increases with input size, measurements were also conducted over the range of 100–1000 vertices as a scalability test beyond the intended input scale.
5.2.2 Boolean Operation Performance
The measurement results are shown in Table 4. In 3D space, for the combinations tested, the computation time was generally on the order of several tens of milliseconds. In contrast, in 4D space the computation time increased, reaching approximately $300\text{\,}\mathrm{ms}$ for Cube–Cube cases and exceeding $2\text{\,}\mathrm{s}$ for Random–Random cases in some instances.
5.3 Rendering Performance
The measurement results are shown in Table 5. Within the measured range, the FPS decreased as the total number of facets increased. Under the experimental conditions used in this study, when the total number of facets was 470 (corresponding to 10 objects), the average FPS was approximately 38.6. These values are recorded as reference values for the present implementation and environment.
5.4 Practical Scope of Application and Usage Scenarios
In this section, based on the measurement results of the 4D implementation, we describe the types of applications that can be realistically supported by the current implementation.
Regarding visualization, in the present experiments, the frame rate decreased as the number of mesh facets increased. This may result in a loss of usability for continuous interaction. Therefore, when using this implementation for exploration or observation with continuous operations, a practical guideline is to keep the total number of facets at approximately 470 or fewer.
For Boolean operations, immediate per-frame updates are not required, and a workflow of “operation → waiting for a certain period → result confirmation → next operation” can be viable. In the experiments, processing time tended to increase as complexity grew, and when it reached the order of several seconds, the pace of trial-and-error interaction was degraded. For this reason, the present implementation is realistically suited to research use cases or prototyping scenarios involving a small number of edits or validations, where waiting for results is acceptable. In contrast, usage scenarios that require continuous immediate responses, such as editing tasks that repeatedly apply Boolean operations at high frequency to large meshes with many facets, may become impractical with this implementation due to increased processing time as the number of facets grows. Accordingly, such scenarios are positioned outside the scope of the implementation and evaluation presented in this paper.
However, as a workflow-level mitigation strategy, a proxy-based approach can be considered as one possible option. In this approach, trial-and-error operations are performed on simplified meshes, and the operations are applied to high-resolution models only at the final stage.
5.5 Discussion and Limitations
The definition of interactivity adopted in this paper holds regardless of computation time; however, practical usability is constrained as computation time increases. The experimental results suggest that when handling meshes with a large number of facets, processing time can reach the order of several seconds, leading to the existence of a practical transition region in which usage patterns that proceed by confirming results after each operation become constrained by waiting time.
It should be emphasized that this transition region does not represent a strict threshold, but rather a practical consideration based on the experimental conditions of this study. Accordingly, this paper does not generalize acceptable response times or practical boundaries as a single numerical value.
The present implementation employs approximate methods based on floating-point arithmetic, and intersection tests as well as inside-outside classification may become unstable depending on degenerate configurations and the choice of tolerances. Therefore, this paper explicitly states the limits of reliability in terms of these dependencies, and does not specify failure conditions as a single numerical boundary.
The experiments in this implementation focus on computational performance and its scalability, and do not include qualitative evaluations such as user studies or structured feedback from domain researchers.
In addition, the visualization-related limitations, namely information loss and dependence on sampling, are described in Section 3.6.1; quantitative evaluation of these limitations is beyond the scope of this paper.
@@ 6 Conclusion and Future Work
In this paper, we proposed a unified architecture for visualization and simulation based on a design targeting $N$-dimensional space. We presented a configuration that integrates multiple processes, including convex hull mesh generation, Boolean operations, coordinate transformations, and hyperplane slicing, into a single software architecture, and adopted the separation of topology and geometry as a design principle. In the 4D implementation, a non-rigid body simulation based on XPBD was implemented within the same architecture, and it was confirmed that updates to mesh vertex coordinates are reflected in cross-section generation and rendering. In addition, an interaction scheme for high-dimensional exploration was designed, and implemented in the 4D implementation.
Furthermore, as one example of data representation for $N$-dimensional meshes, this study designed the Plex data format and defined its specification.
Although the algorithms used in this study are described in a manner independent of the dimensionality $N$, performance evaluation and implementation-level verification are limited to four dimensions.
Future work includes examining the impact of increased computational cost due to the growth of combinatorial complexity in intersection tests and simplicial subdivision for algorithms in five or more dimensions; investigating more numerically robust and physically accurate $N$-dimensional non-rigid physical models; exploring integration using other phenomena, such as fluid simulations, as input examples; and considering GPU implementations for components that are suitable for parallelization among the processes implemented in a CPU-based manner in this paper. In addition, conducting comparative evaluations with existing computational geometry libraries and previously reported methods, and quantitatively characterizing the trade-offs among computation time, memory usage, and robustness to degenerate configurations, also remain as future tasks. Furthermore, conducting qualitative evaluations also remains as future work. Such evaluations include collecting structured feedback from users and domain researchers on representative exploration and editing tasks.
The design and the 4D implementation presented in this study may also serve as candidates for educational and entertainment applications aimed at understanding and manipulating high-dimensional phenomena.
@@ Acknowledgments
We would like to express our sincere gratitude to Mr. Hiroaki Naito of Waseda Osaka High School for his guidance throughout this research. We also wish to express our deepest appreciation to Professor Yasushi Homma of Waseda University for his valuable professional advice on the mathematical expressions related to geometric Algebra used in this paper. Finally, we would like to thank the anonymous reviewers for their constructive comments, which helped to improve this manuscript.